Circular No. 8654

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 A1 (POJMANSKI) Ephemeris from the orbital elements on IAUC 8653 (H = 10, n = 3):

2006TT α_{2000} δ_{2000} Δ r ϵ β Mag.
Jan. 5 21^{h}28.^{m}87 −67^{o}09.^{′}7 1.520 1.164 50.^{o}0 40.^{o}3 11.4
10 21 16.13 −65 14.8 1.484 1.084 46.9 41.5 11.1
15 21 04.83 −63 12.5 1.441 1.003 44.1 43.0 10.8
20 20 54.47 −60 59.3 1.389 0.924 41.6 45.0 10.5
25 20 44.69 −58 29.1 1.328 0.846 39.5 47.8 10.1
30 20 35.38 −55 32.9 1.259 0.773 37.9 51.5  9.7
Feb. 4 20 26.66 −51 57.4 1.180 0.705 36.6 56.5  9.2

COMETS C/2005 W6–W13 (SOHO) Additional near-sun comets have been found on SOHO website images

(cf. IAUC 8650; WX = W. Xu, JR = J. Ruan) — all Kreutz sungrazers except for C/2005 W9 and C/2005 W11 (which both belong to the Meyer group). C/2005 W6 was small, faint, and stellar in appearance in C3 images, but diffuse and barely discernible in C2 images. C/2005 W7 appeared stellar and slightly elongated in C3 images, reaching mag 7.0 at 11.5R_{⊙} on Nov. 19.888 UT; in C2 images, it was relatively large and bright with a short tail that reached a length of 177′′ at 5.8R_{⊙} on Nov. 20.313. C/2005 W8 was small and stellar in appearance in C3 images, reaching mag 7.6 at 11.5R_{⊙} on Nov. 19.888; in C2 images, it appeared stellar and was reasonably bright with no tail. C/2005 W9 and C/2005 W11 were very faint, small, and stellar in appearance. C/2005 W10 was extremely faint and diffuse. C/2005 W12 was very faint, elongated, and diffuse. C/2005 W13 was quite bright and stellar in appearance on C3 images, reaching mag 6.8 at 8.5R_{⊙} on Nov. 21.929; in C2 images it appeared stellar with no tail.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 W6 Nov. 18.446 15^{h}32.^{m}4 −22^{o}45′ C3/2 TS 2005-Y14
C/2005 W7  19.433 15 27.9 −23 29 C3/2 WX 2005-Y14
C/2005 W8  19.447 15 27.5 −23 23 C3/2 JR 2005-Y14
C/2005 W9  19.564 15 38.3 −18 02 C2  HS 2005-Y15
C/2005 W10  19.814 15 39.4 −21 15 C2  TH 2005-Y15
C/2005 W11  19.979 15 39.9 −18 17 C2  HS 2005-Y15
C/2005 W12  20.354 15 41.9 −21 21 C2  QY 2005-Y15
C/2005 W13  21.221 15 34.9 −23 52 C3/2 TS 2005-Y22
2006 January 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green